# Finite Element Operator Network for Solving Parametric PDEs

### v2

@@ 3.4 Further performance metrics
We present the results comparing generalization capability in Figure 9, when the FEONet is trained only with 30 random input samples (green and yellow bar) (which should be distinguished from 30 input-out data pair for the supervised learning). Figure 9 depicts a comparison of train and test errors between the FEONet, trained through an unsupervised approach with 30 and 3000 input samples, and the DeepONet, trained through supervised learning with 30 input-output data pairs. The disparity between the training error and the test error serves as a reliable indicator of how effectively the models generalize the underlying phenomenon. Notably, the FEONet shows superior generalization compared to the DeepONet, even without requiring input-output paired data. Remarkably, utilizing only 30 input samples for training, the FEONet achieves significantly better generalization and exhibits lower test errors than that of the DeepONet.
@@
One of the intriguing aspects is that the theoretical results on the convergence error rate of the FEONet are also observable in the experimental results. Figure 10 shows the relationship between the test error and the number of elements, utilizing both P1 and P2 basis functions. In 1D problems, the convergence rate is approximately 2 for P1 basis functions and 2.9 for P2 basis functions. Meanwhile, in 2D problems, the convergence rate is approximately 2 for P1 basis functions and 2.5 for P2 basis functions. These remarkable trends confirm the theoretical convergence rates observed in the experimental results.
@@ 4 Conclusions
In this paper, we introduce the FEONet, a novel deep learning framework that approximates nonlinear operators in infinite-dimensional Banach spaces using finite element approximations. Our operator learning method for solving PDEs is both simple and remarkably effective, resulting in significant improvements in predictive accuracy, domain flexibility, and data efficiency compared to existing techniques. Furthermore, we demonstrate that the FEONet is capable of learning the solution operator of parametric PDEs, even in the absence of paired input-output training data, and accurately predicting solutions that exhibit singular behavior in thin boundary layers. The FEONet is capable of learning the operator by incorporating not only external forces but also boundary conditions, variable coefficients, and the initial conditions of time-dependent PDEs. We are confident that the FEONet can provide a new and promising approach to simulate and model intricate, nonlinear, and multiscale physical systems, with a wide range of potential applications in science and engineering.
@@ A.1 Approximation error
Note that the constants appearing in the inequalities above may depend on $h>0$. But as we mentioned before, we assumed that $h>0$ is fixed with the sufficiently small finite element error. Then by the universal approximation property , $\inf_{\alpha\in\mathcal{N}_{n}}\|\alpha-\alpha^{*}\|^{2}_{2}\rightarrow 0$ as $n\rightarrow\infty$, which completes the proof.
@@ A.2 Generalization error
Since the class of neural networks under consideration is uniformly bounded and (10) holds, we see that for any $n\in\mathbb{N}$, $\mathcal{G}_{n}$ is $\tilde{b}$-uniformly bounded for some $\tilde{b}>0$. Then the following lemma is the direct consequence of Theorem A.2 within our setting.
@@
$\sup_{\alpha\in\mathcal{N}_{n}}\left|\mathcal{L}^{M}(\alpha)-\mathcal{L}(\alpha)\right|\leq 2R_{n}(\mathcal{G}_{n})+\frac{\delta}{2}.$ (46)
@@
Proof A.4 (Proof of Theorem 2.7).
By Proposition 2.4 and the definition (25), we obtain
$\|\widehat{\alpha}(n)-\widehat{\alpha}(n,M)\|^{2}_{2}$ $\lesssim\|A\widehat{\alpha}(n)-A\widehat{\alpha}(n,M)\|^{2}_{2}\lesssim\left(\|A\widehat{\alpha}(n)-F\|^{2}_{2}+\|A\widehat{\alpha}(n,M)-F\|^{2}_{2}\right)$ (47)
@@
Letting $M\rightarrow\infty$ on (47), we obtain that
@@
which completes the proof.
@@ A.3 Convergence of FEONet
where we have used the fact that $|\phi_{j}|\leq 1$ for all $1\leq j\leq N(h)$ together with Young’s inequality.
@@ B.1 Types and random generation of input functions for FEONet
B.2 Plots on convergence rates
@@
$-\varepsilon\Delta u+\bm{v}\cdot\nabla u$ $=f(x,y),$ $(x,y)\in D=[-1,1]^{2},$ (51)
@@
where $\varepsilon=0.1$ and $\bm{v}=(-1,0)$. The FEONet was used to approximate the solution of these two equations using P1 and P2 nodal basis functions, and the experiments were conducted with varying domain triangulation to have different numbers of elements to observe the convergence rate. As shown in Figure 10, the observed convergence rate of the FEONet shows the convergence rates close to 2 and 3 for P1 and P2 approximation respectively, which are the theoretical results for the classical FEM. Since the precision scale of the neural network we used here is about $10^{-2}$, we cannot observe the exact trends in rates below this level of error. We expect to see the same trend even at lower errors if we use larger scale and more advanced models than the CNN structure.
B.3 Additional plot on generalization error
@@
Figure 11 displays a comparison of the training and test errors between FEONet (w/o labeled data) and DeepONet (with 30 training data pairs) across different settings as performed in Section 3 (Table 1). We used 3000 input samples to train the FEONet and 30 input-output data pairs to train the DeepONet. Even with the 30 data pairs, there occurred a lot of computational costs to gain the data. Although it is predictable that increasing the number of input-output data pairs for the DeepONet would narrow the gap between train and test errors, preparing a large amount of data causes significant computational costs. As depicted in Figure 11, the generalization error of FEONet is notably smaller in comparison to that of DeepONet when trained with 30 data pairs.

### v3

@@ 3.4 Generalization capacity and convergence rate
Figure 9 presents the results comparing the generalization capability of FEONet when trained with only 30 random input samples (green and yellow bars), which should be distinguished from the 30 input-output data pairs used for supervised learning. More specifically, this figure illustrates the generalization capabilities of FEONet in comparison to DeepONet across the three distinct domain settings discussed in Section 3.1: Domain I (circle), Domain II (square with a hole), and Domain III (polygon). Each bar represents the models’ performance when solving Equation (3.1) within these different geometrical domains. Figure 9 compares the training and test errors of FEONet, trained through an unsupervised approach with 30 and 3,000 input samples, against DeepONet, trained through supervised learning with 30 input-output data pairs. The difference between the training and test errors serves as a reliable indicator of the models’ generalization effectiveness. Notably, FEONet demonstrates superior generalization compared to DeepONet, even without requiring input-output paired data. Remarkably, with only 30 input samples for training, FEONet achieves significantly better generalization and exhibits lower test errors than DeepONet. The experimental details can be found in Appendix B.6.
@@
One of the intriguing aspects is that the theoretical results on the convergence error rate of the FEONet are also observable in the experimental results. Figure 10 shows the relationship between the test error and the number of elements, utilizing both P1 and P2 basis functions. In 1D problems, the convergence rate is approximately 2 for P1 basis functions and 2.9 for P2 basis functions. Meanwhile, in 2D problems, the convergence rate is approximately 2 for P1 basis functions and 2.5 for P2 basis functions. These remarkable trends confirm the theoretical convergence rates observed in the experimental results. The experimental details can be found in Appendix B.5.
@@
Remark 3.1.
As we can see from the left in Figure 10, as we increased the number of elements, i.e., on a high-fidelity mesh, the error began to increase again after a certain point, which means that there is an issue regarding the high-dimensional coefficient output of the neural network. To explain this phenomenon, we theoretically investigated the FEONet in our subsequent paper [20] and successfully identified the cause of such phenomenon. More precisely, we discovered that the condition number of the finite element matrix is closely related to the high-fidelity issue of FEONet, which makes total error increase when the degree of freedom is large. Furthermore, this theoretical finding provides a critical clue for resolving the computational issue arising from high-fidelity meshes. By using a preconditioner to define a loss function, we can significantly reduce the total error and the FEONet can effectively make a solution prediction even for the case of a large number of elements. See [20] for details.
@@ 4 Conclusions
In this paper, we introduce the FEONet, a novel deep learning framework that approximates nonlinear operators in infinite-dimensional Banach spaces using finite element approximations. Our operator learning method for solving PDEs is both simple and remarkably effective, resulting in significant improvements in predictive accuracy, domain flexibility, and data efficiency compared to existing techniques. Furthermore, we demonstrate that the FEONet is capable of learning the solution operator of parametric PDEs, even in the absence of paired input-output training data, and accurately predicting solutions that exhibit singular behavior in thin boundary layers. The FEONet is capable of learning the operator by incorporating not only external forces but also boundary conditions, variable coefficients, and the initial conditions of time-dependent PDEs. On the other hand, our method also has some drawbacks that need to be overcome. For example, since the proposed method is not mesh-free and based on domain meshing, we may face challenges in solving high-dimensional PDEs or handling singular problems, such as shock formation (see, e.g. [23]). Note, however, that since our method is based on the classical FEM, we can directly apply numerical analysis techniques to solve these problems, which will be addressed in future work. If these aspects are effectively supplemented, we are confident that the FEONet can provide a new and promising approach to simulate and model intricate, nonlinear, and multiscale physical systems, with a wide range of potential applications in science and engineering.
@@ A.1 Approximation error
Note that the constants appearing in the inequalities above may depend on $h>0$. But as we mentioned before, we assumed that $h>0$ is fixed with the sufficiently small finite element error. Then by the universal approximation property , $\inf_{\alpha\in\mathcal{N}_{n}}\|\alpha-\alpha^{*}\|^{2}_{2}\rightarrow 0$ as $n\rightarrow\infty$, which completes the proof. ∎
@@ A.2 Generalization error
Since the class of neural networks under consideration is uniformly bounded and (2.9) holds, we see that for any $n\in\mathbb{N}$, $\mathcal{G}_{n}$ is $\tilde{b}$-uniformly bounded for some $\tilde{b}>0$. Then the following lemma is the direct consequence of Theorem A.1 within our setting.
@@
$\sup_{\alpha\in\mathcal{N}_{n}}\left|\mathcal{L}^{M}(\alpha)-\mathcal{L}(\alpha)\right|\leq 2R_{n}(\mathcal{G}_{n})+\frac{\delta}{2}.$ (A.1)
@@
Proof of Theorem 2.8.
By Proposition 2.5 and the definition (2.24), we obtain
$\|\widehat{\alpha}(n)-\widehat{\alpha}(n,M)\|^{2}_{2}$ $\lesssim\|A\widehat{\alpha}(n)-A\widehat{\alpha}(n,M)\|^{2}_{2}\lesssim\left(\|A\widehat{\alpha}(n)-F\|^{2}_{2}+\|A\widehat{\alpha}(n,M)-F\|^{2}_{2}\right)$ (A.2)
@@
Letting $M\rightarrow\infty$ on (A.2), we obtain that
@@
which completes the proof. ∎
@@ A.3 Convergence of FEONet
where we have used the fact that $|\phi_{j}|\leq 1$ for all $1\leq j\leq N(h)$ together with Young’s inequality. ∎
@@ B.1 Types and random generation of input functions for FEONet
B.2 Details on hyperparameters
For the problems under consideration in the paper, we used neural networks, which consist of 6 convolutional layers with swish activation followed by a fully connected layer flattening the output. For 1D problems, we used Conv1D, while Conv2D was used for 2D problems. The FEONet was trained with the LBFGS optimizer along with the following hyperparameters.
  • Maximal number of iterations per optimization step: $10$,
  • Termination tolerance on first-order optimality : $10^{-15}$,
  • Termination tolerance on function value/parameter changes: $10^{-15}$,
  • Update history size: $10$.
For both the DeepONet (or POD-DeepONet) and the PIDeepONet, a fully connected neural network with a depth of 3 and a width of 100 was used for the trunk and branch networks. The two models differ only in the loss function, where the training data and the residual of the PDE were used. The PINN also used a fully connected neural network with a depth of 3 and a width of 100. All these models used the tanh activation function and a learning rate of $10^{-4}$ with the commonly used ADAM optimizer. To ensure the sufficient convergence of the training results for the boundary layer problem, we conducted five experiments with the PINN for $5\times 10^{5}$ iterations and five experiments with the PIDeepONet for $10^{4}$ iterations. Note that computing the residual loss of PDE for every input function $f$ in the PIDeepONet requires significant computation time using the automatic differentiation.
B.3 Comparison of computational time
Once the operator-learning ML models (DeepONet, POD-DeepONet, and FEONet) are trained, they can quickly infer solutions for varying PDE parameters. This capability is a key reason why operator networks are becoming increasingly popular in scientific machine learning. In contrast, solving PDEs with traditional FEM for various parameters requires re-solving the PDE each time the parameters change, incurring computational costs. This challenge is further exacerbated when dealing with nonlinear problems, where iterative methods must be applied for each set of PDE parameters, adding to the computation overhead. As mentioned earlier, FEONet is particularly well-suited to address this issue, highlighting a critical difference between traditional FEM and FEONet. To quantify this advantage, let us consider a scenario where 10,000 parametric PDEs with varying parameters need to be solved—a common requirement in computational material database generation. In such cases, operator learning methods like FEONet are significantly faster than traditional FEM. Table 4 illustrates the inference time comparison based on the number of input samples, underscoring the efficiency of models like FEONet in fast-inference applications.
Table 4: Comparison of computational times for FEM, DeepONet, POD-DeepONet and FEONet with different numbers of input samples.
# input samples 1 10 100 1000 10000
FEM 0.012 (s) 0.141 (s) 1.039 (s) 9.979 (s) 99.826 (s)
DeepONet 0.147 (s) 0.123 (s) 0.251 (s) 0.338 (s) 0.746 (s)
POD-DeepONet 0.170 (s) 0.190 (s) 0.191 (s) 0.398 (s) 0.765 (s)
FEONet (Ours) 0.106 (s) 0.221 (s) 0.315 (s) 0.374 (s) 1.529 (s)
B.4 Extension of FEONet to time-dependent PDEs
We use the FEONet to learn an operator that maps $u_{0}(x)=u(t=0,x)$ to $u(\Delta t,x)$ (more precisely, the coefficients of $u(\Delta t,x)$) in Equation (3.6). The other model then predicts the solution at the next time step solution $u(2\Delta t,x)$ as an input $u(\Delta t,x)$), and continues iteratively. To facilitate this, we employed the implicit Euler method to discretize time with $\Delta t=0.01$, and specifically trained the model using the following loss function:
$\mathcal{L}^{M}(\widehat{u}_{h})=\frac{1}{M}\sum^{M}_{m=1}\sum^{N}_{i=1}\bigg{|}\int_{D}\frac{\widehat{u}_{h}(l\Delta t,x)-\widehat{u}_{h}((l-1)\Delta t,x)}{\Delta t}\phi_{i}\mathrm{d}x\\ +\nu\int_{D}\widehat{u}_{h}(l\Delta t,x)_{x}(\phi_{i})_{x}\mathrm{d}x-b\int_{D}\widehat{u}_{h}(l\Delta t,x)(\phi_{i})_{x}\mathrm{d}x\bigg{|}$ (B.3)
for $l=1,2,...,100$ where $\widehat{u}_{h}=\widehat{u}_{h}(l\Delta t,x;\omega_{m})$. This approach could involve training 100 FEONet models over the interval $[0,1]$ by dividing it into 100 segments. However, for efficiency, we divided the interval into 10 segments and modified a model with $u(t=0,x)$ as input and $[u(t=k\Delta t,x)]_{k=1}^{10}$ as output, and then used the output $u(t=10\Delta t,x)$ as input for the next model. Therefore, this approach involved using 10 separate FEONet models.
B.5 Plots on convergence rates
@@
$-\varepsilon\Delta u+\bm{v}\cdot\nabla u$ $=f(x,y),$ $(x,y)\in D=[-1,1]^{2},$ (B.5)
@@
where $\varepsilon=0.1$ and $\bm{v}=(-1,0)$. The FEONet was used to approximate the solution of these two equations using P1 and P2 nodal basis functions, and the experiments were conducted with varying domain triangulation to have different numbers of elements to observe the convergence rate. As shown in Figure 10, the observed convergence rate of the FEONet shows convergence rates close to 2 and 3 for P1 and P2 approximation respectively, which are the theoretical results for the classical FEM. Since the precision scale of the neural network we used here is about $10^{-2}$, we cannot observe the exact trends in rates below this level of error. We expect to see the same trend even at lower errors if we use larger scale and more advanced models than the CNN structure.
B.6 Additional plot on generalization capacity
@@
Figure 11 displays a comparison of the training and test errors between FEONet (w/o labeled data) and DeepONet (with 30 training data pairs) across different settings as performed in Section 3 (Table 1). We used 3000 input samples to train the FEONet and 30 input-output data pairs to train the DeepONet. Even with the 30 data pairs, there are a lot of computational costs to gain the data. Although it is predictable that increasing the number of input-output data pairs for the DeepONet would narrow the gap between train and test errors, preparing a large amount of data causes significant computational costs. As depicted in Figure 11, the generalization error of FEONet is notably smaller in comparison to that of DeepONet when trained with 30 data pairs.
Acknowledgments
Y. Hong was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2021R1A2C1093579) and by the Korea government(MSIT) (RS-2023-00219980). Jae Yong Lee was supported by Institute for Information & Communications Technology Planning & Evaluation (IITP) through the Korea government (MSIT) under Grant No. 2021-0-01341 (Artificial Intelligence Graduate School Program (Chung-Ang University)). S. Ko was supported by National Research Foundation of Korea Grant funded by the Korean Government (RS-2023-00212227). The authors thank the anonymous referees for their helpful comments that improved the quality of the manuscript.